OpenStudy (anonymous):

Use the definition of derivative to find the derivative of f(x) = 3x + 2. Show your work.

OpenStudy (anonymous):

derivative of 3x + 2 = derivative of 3x + derivative of 2 do you know how to find each of them?

OpenStudy (zzr0ck3r):

what is the derivative of 3x? and what is the dervative of 2(or any constant)?

OpenStudy (anonymous):

3 and 0

OpenStudy (zzr0ck3r):

whats 3 + 0?

OpenStudy (anonymous):

there u go o: 3 + 0

OpenStudy (anonymous):

3

OpenStudy (zzr0ck3r):

ok the problem here is that it says use the deffinition, so lim h->0 of (f(x+h)-f(x))/h

OpenStudy (zzr0ck3r):

so lim h->0 (3(x+h)+2-(3x+2))/h

OpenStudy (anonymous):

\[f'(x)=\lim_{h \rightarrow o} \frac{f(x+h)-f(x)}{h}\] f(x+h)=3(x+h)+2 f(x+h)-f(x)=3x+3h+2-3x-2 =3h \[f'(x)=\lim_{h \rightarrow o} \frac{f(x+h)-f(x)}{h}\] =\[\lim_{h \rightarrow 0}\frac{3h}{h}\] = \[\lim_{h \rightarrow 0}3=3\]

OpenStudy (zzr0ck3r):

= lim h->0 (3x+3h+2-3x-2)/h

OpenStudy (zzr0ck3r):

you can do the rest right?

OpenStudy (anonymous):

sorry but i am very lost

OpenStudy (anonymous):

i know how to find the derivative, but i am confused on how to use the definition to explain it

OpenStudy (zzr0ck3r):

do you see that first function that @THE_PROPHET wrote?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

that is the deffinition of the derivative of a function

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

so since your f(x) = 3x+2 (f(x+h)-f(x))/h = (3(x+h)+2 - (3x+2))/h = (3x+3h+2-3x-2)/h = 3h/h = 3

OpenStudy (anonymous):

perfect, thank you so much!!!!